The mass of magnesium oxide resulting from the combustion of 8 g of the element magnesium


1
Expert's answer
2021-05-06T07:38:32-0400

2Mg + O2 = 2MgO

n(Mg) = m/M = 8g/24g/mol = 0.33 mol

According to the equation

n(MgO) = n(Mg) = 0.33 mol

M(MgO) = 24 + 16 = 40 g/mol

m(MgO) = n*M = 0.33 mol*40 g/mol = 13.2 g